OpenStudy (anonymous):

If f(g(x))=x and f(x) = 3x – 4, find g(x) g(x)=(x-4)/3 g(x)=(x+4)/3 g(x)=(x/3)+4 g(x)=(x/3)-4

OpenStudy (anonymous):

@satellite73

OpenStudy (amorfide):

i am going to assume G is the inverse function of F(x) so to find the inverse function let f(x)= y so you get y=3x-4 to get the inverse function you want to make x the subject add 4 to both sides y+4=3x divide by 3 on both sides (y+4)/3=x now let x=g(x) let y=x (x+4)/3=g(x)